OpenStudy (anonymous):

3 square root 25x^3 can anyone answer this?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

\[3\sqrt{25x^3}=15\sqrt{x^3}\]

OpenStudy (anonymous):

do u mean cube root or 3 times the square root?

OpenStudy (anonymous):

that's how the problem is wrote but how dhid yu get that answer steps!

OpenStudy (anonymous):

Are you asking me?

OpenStudy (anonymous):

sure it doesn't matter i just need help understanding how he got that answer

OpenStudy (anonymous):

the problem does look like \[3\sqrt{25x^3}\]?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Okay, 25 is a perfect square so you can pull it out of the square root by taking its square root and multiplying it by what is outside of the square root (in this case 3). So you pull out 25, take its square root (=5), and multiply that by 3 (5*3=15), then you are left with 15*sqrt(x^3)

OpenStudy (anonymous):

thank you soo much,, i understand it now @ aj